OpenStudy (anonymous):

is gravitational acceleration constant for a free falling body??? i mean g = 9.8m/s^2 is constant for a free falling body??

OpenStudy (dls):

yes

OpenStudy (anonymous):

then why the weight of a body varies with respect to height???

OpenStudy (anonymous):

yes, but a lot of height, enough to drastically change the Earth's gravitational influence over the object.

OpenStudy (dls):

\[\LARGE g=\frac{GM}{(1+\frac{H}{R})^2}\approx GM(1-\frac{2H}{R})\] This is the variation with height,which is approximately to the other one if h<<R

OpenStudy (dls):

If you take it to a height comparable to radius of earth..more than 5% more than radius of earth or so.